### v3

@@ 6.3 Stochastic and Deterministic Solutions
In the following, we compare the quality of stochastic solutions with that of deterministic ones. For that purpose, all 1,215 test instances are solved using our stochastic approach and three deterministic approaches assuming best-case, median, or worst-case travel times. We note that we can adjust the AMP to use deterministic travel times $\tau_{i,j}$ for tasks $i$ and $j$ by assuming $P(t_{i,j}=\tau)=1$ when constructing the respective master problem. We thus obtained a total of four solutions for each instance, one for each type of deterministic travel time and one for stochastic travel times. We then sample 1,000 scenarios for each instance by randomly generating travel times for all pairs of parking positions according to the respective empirical distributions. For each instance and each scenario, we analyze the finish times of all tasks for all four solutions. By aggregating these results for each instance and solution over all 1,000 scenarios, we obtain an empirical service level and objective function value (10), which allows us to further assess the performance of deterministic and stochastic solutions. Furthermore, we use the deterministic framework, i.e., constraints (3) and (4) need to be satisfied. For deterministic travel times, this is equivalent to using hard time windows $\mathrm{ES}_{i},\mathrm{LF}_{i}]$ for each task $i$ and disallowing delays. In the following, we refer to the solution of an instance when using stochastic travel times as the stochastic solution of an instance. For best-, median, and worst-case travel times, we analogously refer to best-, median, and worst-case travel time solutions, respectively.
@@
Complexity Travel Times #Feas. #alpha-feas. #LFv-feas. #Stoch.-feas.
Easy Best 266 11 29 11
Median 265 142 243 142
Worst 261 261 261 261
Stochastic 261 261 261 261
Medium Best 243 8 26 8
Median 228 105 192 103
Worst 203 203 203 203
Stochastic 209 209 209 209
Hard Best 222 0 17 0
Median 161 31 98 31
Worst 109 109 109 109
Stochastic 124 124 124 124
All Best 731 19 72 19
Median 654 278 533 276
Worst 573 573 573 573
Stochastic 594 594 594 594
Table 8 summarizes the feasibility of instances and solutions under various travel time assumptions. For each type of travel time, column “#Feas.” contains the number of instances whose master problem (10)–(13) is feasible, assuming the respective travel times. Furthermore, column “#Alpha-feas.” denotes the number of instances whose solutions satisfy the desired service level $\alpha$ for each task. This requirement is equivalent to satisfying constraint (3). Column “#LFe-feas.” serves a similar purpose and contains the number of instances whose solutions guarantee no delays beyond extended latest finish times. This corresponds to satisfying equality (4). Finally, column “#Stoch.-feas.” contains the number of instances whose solutions were feasible with respect to the prescribed requirements regarding both service levels and maximum delays. This is equivalent to fulfilling constraints (3) and (4) simultaneously. It is clear that a decrease in the available workforce increases the frequency with which the prescribed service level is not satisfied or large delays can not be ruled out. Furthermore, only 31 out of 728 best-case travel time solutions satisfy the desired service level for all tasks, while only 25 solutions are feasible with respect to both service level and maximum delay constraints. When median travel times are assumed, around 49% of all solutions guarantee the desired service level, while only 304 out of 649, i.e., around 47% of solutions also fulfill the maximum delay requirement (4). Thus, if deterministic travel times are assumed, it is likely that either tasks are delayed with a probability of more than $1-\alpha$ or large delays of more than 10 minutes can not be prevented, ultimately leading to frequent delays of multiple aircrafts.
@@
Complexity Travel Times Obj Obj-Pen $\mu_{SL}$ $\sigma_{SL}$ $\min_{SL}$
Easy Best 146.74 100.14 89.69% 7.99% 0.00%
Median 25.02 22.97 98.76% 1.32% 32.70%
Worst 3.21 3.21 100.00% 0.00% 100.00%
Stochastic 2.86 2.86 100.00% 0.03% 90.90%
Medium Best 146.11 107.25 91.39% 6.80% 0.20%
Median 33.51 31.86 98.90% 1.33% 28.50%
Worst 18.39 18.39 100.00% 0.00% 100.00%
Stochastic 15.68 15.65 99.96% 0.14% 90.20%
Hard Best 135.29 89.70 89.70% 9.50% 0.20%
Median 69.62 66.79 98.01% 1.81% 26.40%
Worst 68.43 68.43 100.00% 0.00% 100.00%
Stochastic 59.91 59.69 99.82% 0.18% 90.10%
All Best 154.09 109.35 90.29% 7.19% 0.20%
Median 36.14 34.45 98.67% 1.46% 26.40%
Worst 20.99 20.99 100.00% 0.00% 100.00%
Stochastic 18.25 18.23 99.96% 0.13% 90.10%
@@
Table 9 provides an overview of objective function values and service levels. In order to have a stable foundation for comparison, all data visualized in Table 9 is calculated based on the set of instances for which feasible solutions are obtained when assuming worst-case travel times. This amounts to a total of 578 out of 1,215 instances. Recall that all values depicted here are calculated based on a Monte Carlo simulation of travel times. Columns “Obj” and “Obj-Pen” contain the average objective function value of deterministic solutions applied to stochastic travel times, with and without quadratic penalties for time window violations, respectively. Moreover, columns $\mu_{SL}$, $\sigma_{SL}$, and $\min_{SL}$ represent the average service level for all tasks, its standard deviation, and the minimum service level of any task. On average, stochastic solutions return the smallest objective function values, both with and without penalties, for all instance classes. While stochastic solutions’ objective function values are equal to around 3% of the objective function values of best-case travel time solutions (both excluding penalties, respectively) for easy instances, this relation reduces to around 35% for hard instances. Practically speaking, for easy instances, the total safety time buffer accumulated is 30 times larger when using stochastic rather instead of best-case travel time data. Moreover, for every instance class, stochastic solutions guarantee a very high and stable service level, averaging over 99.5% and having a standard deviation of 0.12%. Furthermore, while median travel time solutions provide relatively good objective function values, which are roughly 20% above stochastic solutions, they frequently provide very poor service level bounds for individual tasks, going as low as 21.8% for harder instances. Therefore, if a deterministic model is sought to solve the problem at hand, using median travel times appears to be the most promising approach. Nevertheless, service levels for individual tasks can still be arbitrarily low and clearly lack a lower bound. Such guarantees can only be made with stochastic solutions, which outperform deterministic travel time solutions by all previously mentioned metrics.
@@
Figure 4 shows the distribution of service levels per task. For this purpose, for each type of travel time, we collected all service levels for each individual task for every instance and plotted their empirical distribution. While stochastic travel times guarantee a service level of at least 90% at all times, median and best-case travel time solutions frequently provide significantly lower values, down to 3% in some cases. Moreover, best-case travel time solutions violate the prescribed minimum service level for 25% of tasks, while for median travel time solutions, this holds in around 8% of cases. This emphasizes the unpredictability of minimum service levels when deterministic travel times are assumed, which ultimately has a significant impact on perceived service quality. Hence, explicitly considering stochastic travel times in the problem formulation yields solutions that seldom violate any time windows, resulting in a high and reliable quality of service.
@@
Figure 5 visualizes the occurrence of time window violations, measured in time steps. Recall that for our purposes, each time step is 2 minutes long. As for Figure 4, we collected all potential time window violations and their lengths for each instance and each scenario. Overall, best-case travel time solutions have higher median time window violations, larger 75%-quantiles, and more outliers than stochastic solutions. At the same time, the median time window violations and 75%-quantiles of median and stochastic solutions are almost identical. Nevertheless, the former solutions exhibit far more large delays, reaching up to 11 time steps, i.e., 22 minutes. In total, assuming deterministic travel times usually returns solutions that cause comparably long delays, are at risk of delaying tasks by significant amounts, and lead to highly volatile delays, which are several magnitudes larger than the maximum acceptable delay $\mathrm{LF}_{i}^{\mathrm{e}}-\mathrm{LF}_{i}$. These effects lead to passengers becoming dissatisfied and baggage handling operators having to pay financial penalties. Using stochastic travel times greatly benefits a solution’s quality in these regards, typically causing few, small delays and ruling out undesirably long delays.
To summarize the above findings, we observe that solutions based on deterministic travel times, be they best-case, median, or worst-case, return strategies that do not live up to the set service quality. Explicitly incorporating stochastic travel times into the model allows for solutions that efficiently use the available workforce to increase safety time buffers, guarantee a stable service level, and prevent large delays.
@@ 7 Conclusion
We have examined the problem of forming and routing worker teams, consisting of workers of different hierarchical skill levels. We extended the works of Dall’Olio and Kolisch (2023) by considering stochastic travel times across the apron. For this purpose, we adjusted their MIP formulation to suit stochastic travel times, provide a service level guarantee, and disallow extreme delays for each task. Furthermore, we introduced an alternative master problem formulation abbreviated as DRMP. Moreover, we provided an adjusted dominance rule used to solve the pricing problem associated with said model, allowing us to solve multiple pricing networks at once, at the expense of only a slight increase in runtime. This result ultimately made the usage of said formulation practically feasible. We then combined the DRMP with the formulation proposed by the aforementioned authors to develop a new solution approach based on a Branch-Price-and-Cut approach. More specifically, we switch between two master problem formulations, depending on the characteristics of the current (integer) solution and the position within the branching tree. Furthermore, we introduced an improved branching strategy based on the worst-case finish times of individual tasks, which was able to greatly improve the solution quality and algorithm’s convergence speed. Moreover, we applied an exact separation procedure for rank-1 Chvátal-Gomory cuts and carefully designed the decision process of when and how to separate said cuts. Our computational studies indicate that our proposed algorithm is able to outperform the standard Branch-Price-and-Check approach by Dall’Olio and Kolisch (2023), solving more instances to optimality and significantly reducing optimality gaps. Additionally, we observed that including stochastic travel times into the model is a superior approach to assuming deterministic travel times, as the solutions obtained allow for earlier expected termination of tasks, less frequent and smaller delays, and a guaranteed, stable service level and maximum possible delay.

### v4

@@ 6.3 Stochastic and Deterministic Solutions
In the following, we compare the quality of stochastic solutions with that of deterministic ones. For that purpose, all 1,215 test instances are solved using our stochastic approach with varying workforce quantile levels $\gamma\in\{0.5,0.7,0.9,1\}$ and four deterministic approaches assuming best-case, mean, median, or worst-case travel times. We note that we can adjust the AMP to use deterministic travel times $\tau_{i,j}$ for tasks $i$ and $j$ by assuming $P(t_{i,j}=\tau)=1$ when constructing the respective master problem. Due to (3)–(4), an instance that is feasible for stochastic travel times is not necessarily feasible for deterministic travel times. Similarly, feasibility with respect to best-, mean- or median-case travel times does not imply feasibility with respect to mean-, median- or worst-case travel times, respectively. We thus obtained a total of up to four solutions for each instance, one for each type of deterministic travel time and one for stochastic travel times. Note that, if, for a given instance and a given type of travel time, no solution is obtained, the underlying instance is infeasible for the given (deterministic or stochastic) type of travel time, but it might be feasible for a different type of travel time.
We then sample $N\in\mathbb{N}$ scenarios for each instance by randomly generating travel times for all pairs of parking positions and each time bin according to the respective empirical distributions. For each instance and each scenario, we analyze the finish times of all tasks for all four solutions. By aggregating these results for each instance and solution over all $N$ scenarios, we obtain an empirical service level and objective function value (10), which allows us to further assess the performance of deterministic and stochastic solutions. Furthermore, we impose that constraints (3) and (4) need to be satisfied. For deterministic travel times, this is equivalent to using hard time windows $\mathrm{ES}_{i},\mathrm{LF}_{i}]$ for each task $i$ and disallowing delays. Given a scenario of travel times, a plan (i.e., a solution obtained by solving (6–(9)) is executed as follows. Tours are sorted in ascending order with respect to their scheduled leave time and iteratively implemented. If no sufficient workforce is available to start a planned tour at the scheduled leave time, the tour is postponed until enough workers become available. In the following, we refer to the solution of an instance when using stochastic travel times as the stochastic solution of an instance. For best-, median, and worst-case travel times, we analogously refer to best-, median, and worst-case travel time solutions, respectively. The computation of an appropriate scenario count $N$ is done using Sample Average Approximation techniques. For additional details, we refer to J. In the following, all evaluations are performed based on $500$ scenarios per instance.
@@
Travel Times #Feas. #alpha-feas. #LFe-feas. #Stoch.-feas.
Best 720 30 143 25
Mean 657 282 561 278
Median 657 278 558 273
Worst 582 582 582 582
Stochastic $\gamma=0.5$ 611 608 611 608
Stochastic $\gamma=0.7$ 605 602 605 602
Stochastic $\gamma=0.9$ 601 598 601 598
Stochastic $\gamma=1$ 600 597 600 597
Table 6 summarizes the feasibility of instances and solutions under various travel time assumptions. For stochastic travel times, column Travel Times also denotes the quantile $\gamma$ used for the workforce feasibility constraints (8) and (12). For each type of travel time, column “#Feas.” contains the number of instances for which the master problem (10)–(13) is feasible, assuming the respective travel times. Furthermore, column “#Alpha-feas.” denotes the number of instances for which the solutions satisfy the desired service level $\alpha$ for each task. This requirement is equivalent to satisfying constraint (3). Column “#LFe-feas.” serves a similar purpose and contains the number of instances whose solutions guarantee no delays beyond extended latest finish times. This corresponds to satisfying equality (4). Finally, column “#Stoch.-feas.” contains the number of instances for which the solutions are feasible with respect to the prescribed requirements regarding both service levels and maximum delays. This is equivalent to fulfilling constraints (3) and (4) simultaneously. Only 30 out of 720 best-case travel time solutions satisfy the desired service level for all tasks, while only 25 solutions are feasible with respect to both service level and maximum delay constraints. When median travel times are assumed, around 40% of all solutions guarantee the desired service level and fulfill the maximum delay requirement (4). Thus, if deterministic travel times are assumed, it is likely that either tasks are delayed with a probability of more than $1-\alpha$ or large delays of more than 10 minutes can not be prevented, ultimately leading to frequent delays of multiple aircraft. Moreover, it can be observed that a decreasing value for $\gamma$ only slightly increases the number of feasible instances. Hence, planning workforce availability in constraints (8) and (12) using a quantile $\gamma\in(0,1)$ rather than worst-case travel times does not appear to have a significant impact on instance feasibility. Instead, instance infeasibility appears to be caused mainly by the service level constraints (3) and the presence of latest finish times (4).
@@
Travel Times Obj Obj-Pen $\mu_{SL}$ $\sigma_{SL}$ $\min_{SL}$ VSS EVPI
Best 133.36 97.27 91.78% 6.41 0.00%
Mean 39.81 37.43 98.48% 1.72 9.30%
Median 39.83 37.44 98.46% 1.74 9.30%
Worst 21.53 21.53 99.99% 0.04 89.3%
Stochastic $\gamma=0.5$ 16.99 16.96 99.96% 0.13 87.20% 22.84 4.19
Stochastic $\gamma=0.7$ 17.81 17.78 99.96% 0.12 87.20% 22.01 4.66
Stochastic $\gamma=0.9$ 18.50 18.45 99.95% 0.13 88.70% 21.33 5.31
Stochastic $\gamma=1$ 19.07 19.03 99.95% 0.14 88.70% 20.75 5.57
@@
Table 7 provides an overview of objective function values, service levels, and the value of the stochastic solutions (VSS) as well as the expected value of perfect information (EVPI). In order to have a stable foundation for comparison, all data visualized in Table 7 is calculated based on the set of instances for which feasible solutions are obtained when assuming worst-case travel times. This amounts to a total of 582 out of 1,215 instances. Note that feasibility with respect to worst-case travel times implies feasibility for all (deterministic and stochastic) types of travel times. Recall that all values depicted here are calculated based on a simulation of travel times. Columns “Obj” and “Obj-Pen” contain the average objective function value of deterministic solutions applied to stochastic travel times, with and without quadratic penalties for time window violations, respectively. Moreover, columns $\mu_{SL}$, $\sigma_{SL}$, and $\min_{SL}$ represent the average service level for all tasks, its standard deviation, and the minimum service level of any task. On average, stochastic solutions return the smallest objective function values, both with and without penalties. Furthermore, stochastic solutions’ objective function values are equal to around 15% of the objective function values of best-case travel time solutions (both excluding penalties, respectively), while this relation reduces to around 46% for mean and median travel times. Practically speaking, the total safety time buffer accumulated by stochastic solutions is twice as large when using stochastic rather than mean or median-case travel time data. Column VSS, which contains the average value of the stochastic solution over all instances, indicates a reduction of objective values by up to 22.84 time periods when using stochastic instead of mean travel times during planning. At the same time, the EVPI ranges between 4.19 and 5.57, depending on the value of $\gamma$. Thus, one can conclude that most of the improvement comes from modelling stochasticity rather than from perfect information. Thus, explicitly accounting for stochastic travel times during planning can help realize the majority of the improvement gains from considering uncertainty and greatly improve overall service quality. Moreover, for every instance class, stochastic solutions guarantee a very high and stable service level, averaging over 99.5% and having a standard deviation of 0.12%.
Additionally, the solutions obtained for different levels of $\gamma\in\{0.5,0.7,0.9,1\}$ exhibit several noteworthy properties. First, the average objective value decreases for a decreasing value of $\gamma$. Recall that, in constraint (8), a smaller value of $\gamma$ implies earlier worker availability within the master problem. An earlier worker release allows the inclusion of additional tours in the solution. Although the simulation results show that this leads to more frequent delays, the efficiency gains from executing tasks in separate (additional) tours rather than as part of another tour outweigh the objective value loss from those delays. Consequently, solutions obtained with smaller $\gamma$ values achieve strictly lower expected objective values, even for $\gamma$ as low as $0.5$.
Second, while the minimum service level is close to 90% for $\gamma\in\{0.9,1\}$, this value drops to 87.2% for $\gamma\in\{0.5,0.7\}$. Note that the violation of time windows in the former case can be attributed to the fact that, as elaborated on in Section 2, travel time distributions are propagated based on median finish times at the prior task. However, during simulation, travel times are drawn from distributions based on the actual time bin, leading to potential discrepancies that sometimes slightly reduce the service level. Nevertheless, this effect appears to be rather small. At the same time, for small values of $\gamma$, delays become more frequent, leading to smaller minimum service levels. Hence, while smaller $\gamma$ values indicate an improving performance on average, the prescribed service level is more likely to be violated. Thus, if satisfying the prescribed service level is of utmost importance, one should impose a value $\gamma\geq 0.9$ to minimize the probability of violating said service level.
Furthermore, all deterministic travel times yield very poor service-level bounds for individual tasks, going as low as 9.3% for mean and median travel times. Therefore, if a deterministic model is sought to solve the problem at hand, using mean or median travel times appears to be the most promising approach. Alternatively, worst-case travel times provide better objective values and service levels, yet they frequently fail at finding a feasible solution in the first place. In either case, service levels for individual tasks can still be arbitrarily low and clearly lack a lower bound. Such guarantees can only be made with stochastic solutions, which outperform deterministic travel time solutions by all previously mentioned metrics.
@@
Figure 5 shows the distribution of service levels per task. For this purpose, for each type of travel time, we collected all service levels for each individual task for every instance and plotted their empirical distribution. While stochastic travel times guarantee a service level of at least 90% at almost all times, median and best-case travel time solutions frequently provide significantly lower values, down to 3% in some cases. Moreover, best-case travel time solutions violate the prescribed minimum service level for 22% of tasks, while for median travel time solutions, this holds in around 6% of cases. This emphasizes the unpredictability of minimum service levels when deterministic travel times are assumed, which ultimately has a significant impact on perceived service quality. Hence, explicitly considering stochastic travel times in the problem formulation yields solutions that seldom violate any time windows, resulting in a high and reliable quality of service.
@@
Figure 6 visualizes the occurrence of time window violations, measured in time steps. Recall that for our purposes, each time step is 2 minutes long. As for Figure 5, we collected all potential time window violations and their lengths for each instance and each scenario. Overall, best-case travel time solutions have higher median time window violations, larger 75%-quantiles, and more outliers than stochastic solutions. At the same time, the median time window violations and 75%-quantiles of median and stochastic solutions are almost identical. Nevertheless, the former solutions exhibit far larger delays, reaching up to 12 time steps, i.e., 24 minutes. In total, assuming deterministic travel times usually returns solutions that cause comparably long delays, are at risk of delaying tasks by significant amounts, and lead to highly volatile delays, which are several magnitudes larger than the maximum acceptable delay $\mathrm{LF}_{i}^{\mathrm{e}}-\mathrm{LF}_{i}$. These effects lead to passengers becoming dissatisfied and baggage handling operators having to pay financial penalties. Using stochastic travel times greatly benefits a solution’s quality in these regards, typically causing few, small delays and ruling out undesirably long delays. Moreover, the slight increase in the frequency and magnitude of time window violations for larger values of $\gamma$ can likely be attributed to the structural composition of the solutions: for smaller $\gamma$, tasks that would otherwise appear at the end of long tours - and are therefore particularly delay-prone - are partially shifted to additionally constructed tours, thereby reducing their exposure to cumulative travel time uncertainty.
To summarize the above findings, we observe that solutions based on deterministic travel times, be they best-case, median, or worst-case, return strategies that do not live up to the set service quality. Explicitly incorporating time-dependent stochastic travel times into the model and ensuring workforce feasibility only in a percentage of scenarios allows for solutions that efficiently use the available workforce to increase safety time buffers, guarantee a stable service level, and prevent large delays.
@@ 7 Conclusion
We have examined the problem of forming and routing worker teams, consisting of workers of different hierarchical skill levels. We extended the works of dallolio_temp in several dimensions. From a modelling point of view, we introduce time-dependent travel times across the apron. Moreover, we introduce a quantile-based workforce constraint formulation. Instead of assuming maximum travel times across the apron, we incorporate the $\gamma$-quantile scenario of each travel time distribution. This allows explicit control over the conservatism of workforce planning and can reduce workforce underutilization.
Furthermore, we extended the methodology by dallolio_temp in three aspects. First, we introduced an improved branching strategy based on the worst-case finish times of individual tasks, which was able to greatly improve the solution quality and the algorithm’s convergence speed. Second, we applied an exact separation procedure for rank-1 Chvátal-Gomory cuts as proposed by petersen2008chvatal and carefully designed the decision process of when and how to separate said cuts. Third, we introduce an alternative master problem formulation, abbreviated as DRMP. Based on said formulation and the ARMP formulation by dallolio_temp, we propose a novel solution methodology which we term Branch-Price-Cut-and-Switch. In this context, we switch between two master problem formulations, depending on the characteristics of the current (integer) solution and the position within the branching tree.
Our computational studies indicate that our proposed algorithm is able to outperform the standard Branch-Price-and-Check approach by dallolio_temp as well as the approach by yuan2015branch, solving more instances to optimality and significantly reducing optimality gaps. Additionally, we observed that including stochastic travel times into the model is a superior approach to assuming deterministic travel times, as the solutions obtained allow for earlier expected termination of tasks, less frequent and smaller delays, and a guaranteed, stable service level and maximum possible delay. Most importantly, stochastic planning and the introduction of quantile-based workforce constraints enable the operator to realize the majority of the benefit of modelling stochasticity, as well as to reduce idle times and thus derive efficient solutions.
@@
Acknowledgements
Appendix A List of Symbols
Sets
$\mathcal{B}$ set of time bins (disjoint decomposition of the planning horizon)
$B_{i,j}^{k}$ (finite) support of travel time delays for tasks $i\in\mathcal{I}$ and $j\in\mathcal{I}$ (or tasks and the depot) and time bin $\mathcal{B}_{k}\in\mathcal{B}$
$\mathcal{I}$ Set of tasks
$\mathcal{I}^{r}$ Set of tasks executed by route $r\in\mathcal{R}$